OpenStudy (anonymous):

solve the equation 3x5=2(2x-7) (I then got 3x+5=4x-14. I am stuck because when I continue, the answer doesn't make sense.)

OpenStudy (nincompoop):

3x5? like 15? or is that a variable x raised to 5 like x^5?

OpenStudy (anonymous):

I'm assuming the equation is 3x + 5 = 2(2x-7) as you didn't include the plus sign the first time you wrote the equation. 3x + 5 = 4x - 14 3x + 19 = 4x x = 19

jimthompson5910 (jim_thompson5910):

the goal is to get all the x terms to one side, everything else to the other side 3x+5=4x-14 3x+5+14=4x -14+14 ... Add 14 to both sides. 3x+19 = 4x 19 = 4x-3x ... Subtract 3x from both sides. 19 = 1x 19 = x x = 19

OpenStudy (nincompoop):

i don't know how 3x5 became 3x+5 but in the natural world we can call that a product of sorcery

OpenStudy (anonymous):

@nincompoop She mentioned the equation twice and only forgot the plus the first time.

jimthompson5910 (jim_thompson5910):

probably just a typo and meant to say 3x+5 instead of 3x5

OpenStudy (nincompoop):

or maybe the second one with +5 is a typo too

OpenStudy (anonymous):

thank you every one! sorry about the typo. I forgot to add the + sign after 3x. Thank you very much ^.^

OpenStudy (anonymous):

@nincompoop The answer wouldn't be an integer if it was treated as 3 * 15.

OpenStudy (anonymous):

I meant to say 3 * 5. Now I'm making typos too.

OpenStudy (nincompoop):

contagious